OpenStudy (anonymous):

for a 50,000 fundrasing campaign, 98% of the funds have been raised. Which of the three forms would ypu use to mentally caculate tgis amount?

OpenStudy (anonymous):

I'd multiply mentally by decimals. .98 * 50,000 .98 * 50 = 49 *1000 = 49,000 So decimals. Probably not fractions or percents.